Partially dried clay mineral contains 8% water.The original sample contained 12% water and 45% silica.The % of silica in partially dried sample is nearly ___?

Grade:11

2 Answers

Kaushik
10 years ago

Let mass of original clay sample taken be 100g.

clearly,12g of water and 45 g of silica are present in the sample.

Let mass of evaporated water be x g.

(12-x)/(100-x) * 100=8

x=100/23 g

% silica in dried sample=45/(100-x) *100

=43.043%

Kaushik
10 years ago

I am so sorry.I did a mistake in calculation in the last step.

Answer is 47%